I need help interpreting this question: two consecutive even integers are such that three times the smaller plus seven times the larger is -166. find the smaller integer.
Help would be appreciated.

Answer by rfer(16322)

3x+7(x+2)=-166
3x+7x+14=-166
10x=-180
x=-180/10
x=-18
x+2=-16